OpenStudy (jesuslover11):

Find the slope of the line through (-9, -10) and (-2, -5). A) - 5/7 B) 7/5 C)5/7 D)- 7/5

OpenStudy (jesuslover11):

Please help!!

OpenStudy (alexandervonhumboldt2):

sloe formula: (y1-y2)/(x1-x2)

OpenStudy (alexandervonhumboldt2):

*slope

OpenStudy (alexandervonhumboldt2):

(-10+5)/(-9+2)=?

OpenStudy (alexandervonhumboldt2):

solve

OpenStudy (alexandervonhumboldt2):

\[\frac{ y_1-y_2 }{ x_1-x_2 }=\frac{ -10-(-5) }{ -9-(-2) }=\frac{ -10+5 }{ -9+2 }=?\]

OpenStudy (michele_laino):

If I call with \((x_1,y_1)=(-9,-10)\), and with \((x_2,y_2)=(-2,-5)\), then the slope \(m\) of such line is given by the subsequent formula: \[\Large m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 5 - \left( { - 10} \right)}}{{ - 2 - \left( { - 9} \right)}} = ...?\]

OpenStudy (jesuslover11):

-5/7?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

-10+5=-5 -9+2=-7

OpenStudy (alexandervonhumboldt2):

thus (-5)/(-7)=?

OpenStudy (michele_laino):

hint: \(-5-(-10)=-5+10=...?\)

OpenStudy (michele_laino):

similarly: \(-2-(-9)=-2+9=...?\)

OpenStudy (jesuslover11):

5/7?

OpenStudy (michele_laino):

that's right!